How much money does a typical family of four spend at a McDonald’s restaurant per visit? The amount is a normally distributed random variable with a mean of $16.40 and a standard deviation of $2.75.

a. Find the probability that a family of four spends less than $10.

b. What is the amount below which only 10% of families of four spend at McDonald’s?


1
Expert's answer
2020-11-26T18:08:30-0500

A typical family of four spends $16.40.

a. z=(x-"\mu")/"\sigma"=(10-16.4)/2.75=-2.33

P(x<z)=0.00990

b. P(x<z)=0.10

z=-1.28

z=(x-"\mu")/"\sigma"

x=z*"\sigma"+"\mu"=-1.28*2.75+16.4=12.88